# Winning Solutions for the Rayan AI Contest: Compositional Retrieval, Zero-Shot Anomaly Detection, and Backdoor Detection

### v1

@@ Abstract
This report presents solutions to three machine learning challenges: compositional image retrieval, zero-shot anomaly detection, and backdoored model detection. In compositional image retrieval, we developed a system that processes visual and textual inputs to retrieve relevant images, achieving 95.38% accuracy and ranking first with a clear margin over the second team. For zero-shot anomaly detection, we designed a model that identifies and localizes anomalies in images without prior exposure to abnormal examples, securing 1st place with 73.14% accuracy. In the backdoored model detection task, we proposed a method to detect hidden backdoor triggers in neural networks, reaching an accuracy of 78%, which placed our approach in second place. These results demonstrate the effectiveness of our methods in addressing key challenges related to retrieval, anomaly detection, and model security, with implications for real-world applications in industries such as healthcare, manufacturing, and cybersecurity. Code for all solutions is available online. [^1] https://github.com/safinal/rayan-ai-contest-solutions
@@ 1 Introduction
This report covers three distinct tasks: compositional image retrieval, zero-shot anomaly detection, and backdoored model detection. The first task involves developing a system capable of retrieving images based on multi-modal queries, combining visual content with textual modifications. The second task focuses on detecting anomalies, particularly in industrial and medical settings. The third task aims to detect backdoor attacks in deep learning models, where a model’s performance is manipulated by malicious inputs or model weights during training. Each challenge presents unique objectives, including efficient retrieval, generalization to unseen distributions, and identifying security vulnerabilities in models.
The evaluation of these solutions is based on a set of standard metrics, such as accuracy, and F1 score, to assess both the quality and robustness of the approaches. In this report, we describe the methodologies we employed, the experiments conducted, and the results obtained, demonstrating the effectiveness of our solutions. Our team, No Trust Issues Here, achieved first place in the competition, with a final score of 95.38% for the compositional image retrieval challenge (Q5), 73.14% for zero-shot anomaly detection (Q6), and 78% for backdoored model detection (Q7). The final leaderboard is shown in Table 1.
Table 1: Final leaderboard showing the top 10 teams and their scores for the three challenges.
Rank Team Q5 (%) Q6 (%) Q7 (%)
1 No Trust Issues Here (Our Team) 95.38 73.14 78
2 AI Guardians of Trust 88.59 66.29 72
3 AIUoK 78.32 72.98 70
4 red_serotonin 85.45 63.51 65
5 Persistence 80.63 61.62 74
6 GGWP 80.97 62.25 63
7 Tempest 69.86 70.88 63
8 AlphaQ 82.89 62.15 60
9 Cogniverse 49.71 61.53 63
10 AUTs - 60.84 79
2 Q5: Compositional Image Retrieval
@@ 2.1.1 Token Classification for Semantic Parsing
Figure 2: Overview of our token classification module.
@@ 2.1.2 Compositional Embedding Arithmetic for Retrieval
The second stage retrieves the target image by modifying the embedding of the reference image according to the objects identified in the query. We evaluated many models from OpenCLIP (ilharco_gabriel_2021_5143773) and selected ViTamin-L-384 (chen2024vitamin) due to its superior performance in preliminary experiments. We fine-tune only the visual head (model.feature_extractor.visual.head), keeping the rest of the backbone fixed to remain within the resource limits.
We optimize the retrieval model using InfoNCE loss (oord2018representation). The objective is to make the transformed embedding of the query image
$\mathbf{v}_{\text{img}}+\mathbf{t}_{\text{pos}}-\mathbf{t}_{\text{neg}}$
similar to the embedding of the ground-truth target image. In practice, the modification embeddings are computed by encoding the text prompts “a photo of a [object]” using the model’s text encoder. During training, this encourages the model to learn a consistent direction in the embedding space for each object addition or removal.
Training uses AdamW (learning rate 0.0001, weight decay 0.01), five training epochs, and a loss temperature of 0.07. We further apply a CosineAnnealingWarmRestarts scheduler with $T_{0}=5$ and $T_{\text{mult}}=2$. Because the dataset contains repeated occurrences of the same target image, we implement a custom sampler to ensure that each batch contains unique target images, which is required for the InfoNCE formulation.
During inference, the Token Classifier identifies all positive and negative objects in the text. We then encode the corresponding text prompts to obtain the positive embeddings $\{\mathbf{t}^{(i)}_{\text{pos}}\}$ and negative embeddings $\{\mathbf{t}^{(j)}_{\text{neg}}\}$. Given the reference image embedding $\mathbf{v}_{\text{img}}$, we compute:
$\mathbf{v}_{\text{target}}=\mathbf{v}_{\text{img}}+\sum_{i}\mathbf{t}^{(i)}_{\text{pos}}-\sum_{j}\mathbf{t}^{(j)}_{\text{neg}}$.
We then compute cosine similarity between $\mathbf{v}_{\text{target}}$ and all database image embeddings, and return the image with the highest similarity score.
@@
Figure 3: Overview of the proposed compositional image retrieval method. The system parses textual modifications using a lightweight token classifier and applies embedding arithmetic to adjust the reference image representation before retrieval.
@@ 2.2 Experimental Results
3 Q6: Zero-Shot Anomaly Detection
@@
(a) Capsule (normal)
(b) Capsule (anomaly)
(c) Capsule (mask)
(d) Photovoltaic (normal)
(e) Photovoltaic (anomaly)
(f) Photovoltaic (mask)
(g) Pill (normal)
(h) Pill (anomaly)
(i) Pill (mask)
@@ 3.2 Experimental Results
Table 3 shows the leaderboard for this problem. Our team, No Trust Issues Here, achieved the highest performance in the challenge. By leveraging the logic detailed in the methodology section, our solution attained a final accuracy of 73.14%, securing the 1st rank on the leaderboard. This performance demonstrated superior generalization across the diverse industrial and medical classes provided in the blind test set.
@@
10 RANA 61.60
4 Q7: Backdoored Model Detection
@@ 4.1.1 Activation Optimization
We use hyperparameters that were stable across datasets: 100 optimization steps, batch size 30, an initial learning rate of 0.1 decayed to 0.001, and 10 warmup steps.
@@ 4.1.2 Statistical Outlier Detection
  • Distribution Fitting: We collect the maximal confidence scores across classes, exclude the maximum, and fit the remaining scores to an exponential distribution. We tested all SciPy candidate distributions and found the exponential distribution produced the best separation between clean and backdoored models.
@@ 5 Conclusion
This report addressed three critical challenges in machine learning: compositional image retrieval, zero-shot anomaly detection, and backdoored model detection. In the compositional retrieval task, we developed a system that successfully combined visual and textual inputs to retrieve relevant images, achieving first place with a top accuracy of 95.38%. For zero-shot anomaly detection, we implemented a method that detected and localized anomalies in unseen data, securing first place with an accuracy of 73.14%. In the backdoored model detection task, our approach identified backdoor attacks in deep neural networks, achieving a 78% accuracy, placing second in the competition. These results demonstrate the effectiveness of our methodologies in addressing each problem’s specific challenges.

### v2

# No Trust Issues Here: A Technical Report on the Winning Solutions for the Rayan AI Contest

@@ Abstract
This report presents solutions to three machine learning challenges developed as part of the Rayan AI Contest: compositional image retrieval, zero-shot anomaly detection, and backdoored model detection. In compositional image retrieval, we developed a system that processes visual and textual inputs to retrieve relevant images, achieving 95.38% accuracy and ranking first with a clear margin over the second team. For zero-shot anomaly detection, we designed a model that identifies and localizes anomalies in images without prior exposure to abnormal examples, securing second place with a 73.14% score. In the backdoored model detection task, we proposed a method to detect hidden backdoor triggers in neural networks, reaching an accuracy of 78%, which placed our approach in second place. These results demonstrate the effectiveness of our methods in addressing key challenges related to retrieval, anomaly detection, and model security, with implications for real-world applications in industries such as healthcare, manufacturing, and cybersecurity. Code for all solutions is available online. [^1] https://github.com/safinal/rayan-ai-contest-solutions
@@ 1 Introduction
The solutions detailed in this report were developed for the Rayan AI Contest [^2] https://ai.rayan.global, an international competition hosted by Sharif University of Technology focused on Trustworthy AI. The contest attracted significant global participation, with over 4,600 contestants from more than 60 countries forming 730 teams during the screening phase. From this pool, 100 teams were invited to the main phase, and ultimately, more than 40 teams successfully participated with at least one valid submission. To address the competition’s objectives, we focused on three distinct tasks: compositional image retrieval, zero-shot anomaly detection, and backdoored model detection.
The first task involves developing a system capable of retrieving images based on multi-modal queries. The second focuses on detecting anomalies in industrial and medical settings, and the third aims to identify security vulnerabilities, specifically backdoor attacks, in deep learning models.
The evaluation of these solutions is based on a set of standard metrics, such as accuracy and F1 score, to assess both the quality and robustness of the approaches. Our team, No Trust Issues Here, secured the 1st place overall by achieving state-of-the-art performance across all three tracks. Specifically, we ranked 1st in compositional image retrieval with a score of 95.38%, 2nd in zero-shot anomaly detection with 73.14%, and 2nd in backdoored model detection with 78%. The final leaderboard ranking was determined by Total Points, calculated as:
$\text{Total Points}=\sum_{i=1}^{3}(101-\text{rank}_{i})$
where $\text{rank}_{i}$ denotes the team’s rank in the $i$-th challenge. Our cumulative 298 points placed us at the top of the final leaderboard, as shown in Table 1.
Table 1: Final leaderboard showing the top 10 teams.
Rank Team Rank Total Points
Compositional Image Retrieval Zero-Shot Anomaly Detection Backdoored Model Detection
1 No Trust Issues Here (Our Team) 1 2 2 298
2 Pileh 4 1 6 292
3 AI Guardians of Trust 2 6 4 291
4 AIUoK 9 3 5 286
5 red_serotonin 3 7 9 284
6 Persistence 7 11 3 282
7 GGWP 6 9 11 277
8 Tempest 11 5 13 274
9 AlphaQ 5 10 16 272
10 Cogniverse 20 13 12 258
2 Compositional Image Retrieval
@@ 2.1.1 Token Classification for Semantic Parsing
Figure 2: Overview of the token classification module. (a) Dataset curation using templated queries instantiated with LLM-generated object vocabulary. (b) Fine-tuned DistilBERT model for token-level classification into positive, negative, and other categories, using the curated dataset.
@@ 2.1.2 Compositional Embedding Arithmetic for Retrieval
The second stage retrieves the target image by modifying the vector representation of the query image according to the objects identified in the query text. We evaluated multiple models from OpenCLIP (ilharco_gabriel_2021_5143773) and selected ViTamin-L-384 (chen2024vitamin) due to its superior performance in preliminary experiments. To remain within resource limits, we fine-tune only the visual head (model.feature_extractor.visual.head) while keeping the rest of the backbone fixed.
We optimize the retrieval model using the InfoNCE loss (oord2018representation). The objective is to align the composed query vector with the vector of the ground-truth target image. Let $\mathbf{v}_{\text{query}}\in\mathbb{R}^{d}$ be the query image embedding and $\mathbf{t}_{\text{pos}},\mathbf{t}_{\text{neg}}\in\mathbb{R}^{d}$ be the embeddings of the objects to be added and removed, respectively. In practice, the modification vectors are computed by encoding the text prompts “a photo of a [object]” using the model’s text encoder. During training, this encourages the model to learn a consistent direction in the embedding space for each object addition or removal.
Training utilizes the AdamW optimizer (learning rate $10^{-4}$, weight decay $0.01$) over five epochs with a loss temperature of $0.07$. We employ a CosineAnnealingWarmRestarts scheduler with $T_{0}=5$ and $T_{\text{mult}}=2$. Because the dataset contains repeated occurrences of the same target image, we implement a custom sampler to ensure that each batch contains unique target images, which is required for the InfoNCE formulation.
During inference, the Token Classifier identifies the sets of positive objects $\mathcal{P}$ and negative objects $\mathcal{N}$ from the query text. We encode these to obtain the corresponding embedding vectors $\mathbf{t}_{\text{pos}}^{(i)}$ and $\mathbf{t}_{\text{neg}}^{(j)}$. Given the query image vector $\mathbf{v}_{\text{query}}$, the predicted target vector $\mathbf{v}_{\text{target}}$ is computed as:
$\mathbf{v}_{\text{target}}=\mathbf{v}_{\text{query}}+\sum_{i\in\mathcal{P}}\mathbf{t}_{\text{pos}}^{(i)}-\sum_{j\in\mathcal{N}}\mathbf{t}_{\text{neg}}^{(j)}$
The system then computes the cosine similarity between $\mathbf{v}_{\text{target}}$ and all image vectors in the database, returning the image with the highest similarity score.
@@
Figure 3: Overview of the proposed compositional image retrieval method. The system parses textual modifications using our fine-tuned token classifier and applies embedding arithmetic to adjust the reference image representation before retrieval.
@@ 2.2 Experimental Results
3 Zero-Shot Anomaly Detection
@@ 3.2 Experimental Results
Table 3 shows the leaderboard for this problem. Our team, No Trust Issues Here, achieved the second highest performance in the challenge. By leveraging the logic detailed in the methodology section, our solution attained a final score of 73.14%, securing the 2nd rank on the leaderboard. This performance demonstrated superior generalization across the diverse industrial and medical classes provided in the blind test set.
@@
4 Backdoored Model Detection
@@ 4.1.1 Activation Optimization
We use hyperparameters that were stable across datasets: 100 optimization steps, batch size 30, an initial learning rate of $10^{-1}$ decayed to $10^{-3}$, and 10 warmup steps.
@@ 4.1.2 Statistical Outlier Detection
  • Distribution Fitting: We collect the maximal confidence scores across classes, exclude the maximum, and fit the remaining scores to an exponential distribution. We tested all SciPy (2020SciPy-NMeth) candidate distributions and found the exponential distribution produced the best separation between clean and backdoored models.
@@ 5 Conclusion
This report addressed three critical challenges in machine learning: compositional image retrieval, zero-shot anomaly detection, and backdoored model detection. In the compositional retrieval task, we developed a system that successfully combined visual and textual inputs to retrieve relevant images, achieving first place with a top accuracy of 95.38%. For zero-shot anomaly detection, we implemented a method that detected and localized anomalies in unseen data, securing second place with an accuracy of 73.14%. In the backdoored model detection task, our approach identified backdoor attacks in deep neural networks, achieving a 78% accuracy, placing second in the competition. These results demonstrate the effectiveness of our methodologies in addressing each problem’s specific challenges.